I don't understand the wording or what formula I need to plug this problem into. Find the equations of the circles that are tangent to the x axis and have a radius of length five units. In each case, the abscissa of the center is -3.(There is more than one circle that satisfies these conditions)
Answer by venugopalramana(3286)

Find the equations of the circles that are tangent to the x axis and have a radius of length five units. In each case, the abscissa of the center is -3.(There is more than one circle that satisfies these conditions)
WHEN THE CIRCLE IS TANGENT TO X AXIS ,IT MEANS ITS CENTRE IS AT A DISTANCE EQUAL TO RADIUS OF THE CIRCLE FROM THE X AXIS .THAT IS Y COORDINATE OF CENTRE OF CIRCLE = 5 OR -5
ABSCISSA OF CENTRE IS GIVEN AS -3.
HENCE CENTRE IS EITHER (-3,5).....OR (-3,-5)..ITS RADIUS =5
HENCE EQN OF 2 POSSIBLE CIRCLES ARE
(X+3)^2+(Y-5)^2 = 5^2=25...................I
OR
(X+3)^2+(Y+5)^2=25........................II